# Ikeda type lift on ${\rm SO}(3,n+1)$

### v2

@@ 4. Eisenstein series
In this section, we follow [23, Section 4]. For any $l>n+1$ and a Schwartz-function $\Phi_{\mathbb{f}}$ on $V(\mathbb{A}_{\bf f})$, we define a section $f$ of ${\rm Ind}^{G(\mathbb{A})}_{P(\mathbb{A})}\,|\nu|^{s-\frac{n+2}{2}}$ (normalized induction) by
@@
$f_{{\rm fte}}(g_{\mathbb{f}},\Phi_{\mathbb{f}},s)=\int_{{\rm GL}_{1}(\mathbb{A}_{\mathbb{f}})}|t|^{s}\Phi_{\mathbb{f}}(tg_{\mathbb{f}}e)dt,\ g_{\mathbb{f}}\in G(\mathbb{A}_{\mathbb{f}})$
and $f_{l,\infty}(g_{\infty},s)$ is the $V_{l}={\rm Sym}^{2l}(\mathbb{C}^{2})$-valued, $K_{\infty}$-equivalent section of ${\rm Ind}^{G(\mathbb{R})}_{P(\mathbb{R})}\,|\nu|^{s-\frac{n+2}{2}}$. See Section 4.2 for a precise definition. Then, we define the associated Eisenstein series as
@@
where $w\in G={\rm SO}(V)$ is defined by $w\cdot e=f$, $w\cdot f=e$, and $w|_{V^{\prime}}={\rm id}_{V^{\prime}}$. Notice that $J(s,\eta,\Phi_{p})$ is also defined for $\eta\in V^{\prime}(\mathbb{Q}_{p})$ with $q(\eta)\neq 0$. Further, by [23, p.633, line -6], $J(s,\eta,\Phi_{p})=\displaystyle\int_{{\rm GL}_{1}(\mathbb{Q}_{p})}\int_{V^{\prime}(\mathbb{Q}_{p})}\psi_{p}((\eta,x))\Phi_{p}(t,tx,-tq^{\prime}(x)))|t|^{s}_{p}\,dxdt$. Since $t$ runs over $\mathbb{Z}^{\times}_{p}$, $\Phi_{p}(t,tx,-tq^{\prime}(x)))$ is preserved under the transformation $x\mapsto x+x_{0}$ for any $x_{0}\in V^{\prime}(\mathbb{Z}_{p})$. By the change of variables, $J(s,\eta,\Phi_{p})=\psi((\eta,x_{0}))J(s,\eta,\Phi_{p})$ for any $x_{0}\in V^{\prime}(\mathbb{Z}_{p})$. Thus, $J(s,\eta,\Phi_{p})=0$ if $\eta\not\in V^{\prime}(\mathbb{Z}_{p})$.
The following lemma is an analogue of [16, p.591, Lemma 3.4] and it is proved by the same argument therein:
@@
Keep the notation as above. For each $g_{p}={\rm diag}(t_{p},m_{p},t_{p}^{-1})n_{p}\in P(\mathbb{Q}_{p})=M(\mathbb{Q}_{p})N(\mathbb{Q}_{p})$, it holds
$J(s,\eta,g_{p}\cdot\Phi_{p})=|t_{p}|^{\frac{n+2}{2}}_{p}J(s,g_{p}^{-1}\eta,\Phi_{p}).$
@@ 4.1. $p$-adic computation: Siegel series
$B_{r,\eta}=\int_{p^{-r}V^{\prime}(\mathbb{Z}_{p})}\psi((\eta,x))\text{char}(p^{r}q(x)\in\mathbb{Z})\,dx=p^{r(n+2)}\int_{V^{\prime}(\mathbb{Z}_{p})\atop q(x)\equiv 0\,\text{mod $p^{r}$}}\psi\Big(\frac{(\eta,x)}{p^{r}}\Big)\,dx.$
@@ 5. The Weil representation of the Jacobi group $J$
We refer to [9] for the basics of the Weil representation. Let $\psi:\mathbb{Q}\backslash\mathbb{A}\longrightarrow\mathbb{C}^{\times}$ be the standard additive character (cf [15, Section 4.2]). For each $S\in\mathbb{Q}^{\times}$, let $\psi_{S}=\psi(S\ast)=\otimes^{\prime}_{p}\psi_{S,p}$ where we also regard it with an additive character on $Z(\mathbb{A})$ via the identity $Z(\mathbb{A})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{A},\ v(0,0,z)\mapsto z$.
@@ 6.2. Fourier-Jacobi expansions of Eisenstein series
Let $\iota_{{\rm SL}_{2}}=\left(\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right)$ be the Weyl element of ${\rm SL}_{2}$. The following theorem is proved by a standard way (see the proof of [9, p.627, Theorem 3.2]) with the double coset decomposition in Section 6.1. Thus, we omit the proof.
@@ 7.1. Preliminaries at the archimedean place
$\lambda=\sqrt{\beta}\sqrt{2\pi|S|},\ \mu=2\pi\beta\Big(|a|+|S|t^{2}_{n}+|S|(x^{\prime},x^{\prime})\Big)$
@@ 7.2. Computation of Fourier-Jacobi coefficients
Note that (7.16) is a finite sum (see [15, Section 5.2]).
@@
$n(-\alpha,\alpha x_{\infty},\frac{\alpha}{2}\sigma(x_{\infty},x_{\infty}))v(x_{\infty},0,0){\rm diag}(\sqrt{\beta},\sqrt{\beta}^{-1})\in N(\mathbb{R})M(\mathbb{R})$
@@
$\phantom{xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx}=\varepsilon^{n}_{\infty}\frac{\gamma_{\infty}(1)}{\gamma_{\infty}(S^{n})}\beta^{\frac{n}{4}}\Phi_{\infty}(\sqrt{\beta}x_{\infty})\psi_{\infty}(\tfrac{S\alpha}{2}\sigma(x_{\infty},x_{\infty})).$
@@
(this follows from the argument for [15, Section 5.2] and the proof of [16, Lemma 5.4]).
@@ 8. A uniform proof for the main theorems
In this section, we prove Theorem 1.1 and 1.2. We follow the method in [10], [11], [15] rather than representation theoretic arguments [14], [16].
@@ 8.1. A formal series
Recall the Laurent polynomial $\widetilde{Q}_{\eta,p}(X_{p})$ depending on $\eta\in V^{\prime}(\mathbb{Q}_{p})$ and $\Phi_{p}={\rm char}_{V^{\prime}(\mathbb{Z}_{p})}$. We write it as $\widetilde{Q}_{\eta,\Phi_{p}}(X_{p}):=\widetilde{Q}_{\eta,p}(X_{p})$ to indicate independence. By Lemma 4.1 and the computation in Section 4, for each $g_{p}={\rm diag}(t_{p},m_{p},t_{p}^{-1})n_{p}\in P(\mathbb{Q}_{p})=M(\mathbb{Q}_{p})N(\mathbb{Q}_{p})$, we see that
(8.1) $\widetilde{Q}_{\eta,g^{-1}_{p}\Phi_{p}}(X_{p})=|t_{p}|^{\frac{n-2}{2}}_{p}\widetilde{Q}_{g_{p}^{-1}\eta,\Phi_{p}}(X_{p})$
@@ 8.2. A compatible family of Eisenstein series
makes up a compatible family of Eisenstein series in the sense of [11, Section 9] for $n$ even and of [10, Section 10] for $n$ odd.
The case when $S>0$ is similarly handled and we also have a compatible family of Eisenstein series for $v=0,-l^{\prime}$. In fact, we have a family of anti-holomorphic Eisenstein series but the above mentioned results also holds in this setting.
@@ 8.3. Automorphy of the formal series
$F_{f}(g):=F(g;\{\alpha_{p}\}_{p})=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}A(\eta;\{\alpha_{p}\}_{p})(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\quad g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A}),$
@@
Now assume $n$ is odd. Let $D$ be a square-free fundamental discriminant. Then for $\eta=(D,D,\overbrace{0,\ldots,0}^{n},\pm 1,\pm 1)$, $q(\eta)=\pm D$, where we choose a sign so that $\pm D>0$. Then by Theorem 4.3, Theorem 4.9, and (8.3), $A(\eta;\{\alpha_{p}\}_{p})=c(|D|)$. By Kohnen-Zagier [18, Theorem 1], $c(|D|)$ is proportional to the central value of $f\otimes\chi_{D}$. The claim follows from [22, p.654]. ∎
@@
Let $\gamma={\rm diag}(c,m,c^{-1})n\in P(\mathbb{Q})=M(\mathbb{Q})N(\mathbb{Q})$. By (8.1) and (3.2), we have
$F_{f}(\gamma g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}A(\eta;\{\alpha_{p}\}_{p})(\gamma g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(\gamma g_{\infty})$
$\phantom{xxxxx}=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}|c|^{l-\frac{n}{2}}_{\mathbb{A}_{f}}\cdot|c|^{\frac{n+2}{2}}_{\mathbb{A}_{f}}A(\gamma^{-1}\eta;\{\alpha_{p}\}_{p})(g_{\mathbb{f}})|c|^{(l+1)}_{\infty}\mathcal{W}_{2\pi\gamma^{-1}\eta}(g_{\infty}).$
Clearly, $|c|^{l-\frac{n}{2}}_{\mathbb{A}_{f}}\cdot|c|^{\frac{n+2}{2}}_{\mathbb{A}_{f}}|c|^{(l+1)}_{\infty}=|c|^{l+1}_{\mathbb{A}}=1$ since $c\in\mathbb{Q}^{\times}$. Replacing $\gamma^{-1}\eta$ with $\eta$ does not change the conditions in the summation. Thus, we have the claim. ∎
@@
where $\Phi$ runs over $\mathcal{S}(X(\mathbb{A}))$. Notice $H^{{\rm ss}}(\mathbb{Q})$ splits under the double covering $\widetilde{H^{{\rm ss}}(\mathbb{A})}\longrightarrow H^{{\rm ss}}(\mathbb{A})$ and $\Theta_{\psi_{S}}(\widetilde{h}v)$ is left $H^{{\rm ss}}(\mathbb{Q})$-invariant. Then, by the argument for [16, Lemma 5.4], for any fixed $v$ ($l\leq v\leq l$), we may check the $H^{{\rm ss}}(\mathbb{Q})$-invariance of the $v$-th component $((F_{f})_{S,\Phi})_{v}$ (defined as (7.19)) of $(F_{f})_{S,\Phi}$ for any $\Phi\in\mathcal{S}(X(\mathbb{A}))$.
@@
Since $F_{f}$ is right G(^)G($\widehat{\mathbb{Z}}$) by construction and left $P(\mathbb{Q})$-invariant by Lemma 8.3, we may assume $\Phi_{\mathbb{f}}={\rm char}_{V^{\prime}(\widehat{\mathbb{Z}})}$. Assume $S<0$. First, we observe the $l$-th component $((F_{f})_{\xi,S,\Phi})_{l}$ for $\Phi$ whose archimedean component $\Phi_{\infty}$ is Gaussian $\Phi_{S,0}$. Then, Applying Section 8.2 and [11, Theorem 10.2] for $n$ even and [10, Theorem 10.5] for $n$ odd, the $l$-th component $((F_{f})_{\xi,S,\Phi})_{l}$ is automorphic, hence left $H^{\rm ss}(\mathbb{Q})$-invariant. It is also corresponding to a holomorphic modular form of weight $l-\frac{n}{2}+1$. Further, $C_{4}(S,\Phi_{\infty})=C_{2}(S,\Phi_{S,0})$ is non-zero by Remark 7.1. Thus, the formal Fourier-Jacobi coefficient
@@
Now we turn to look the 0-th component $((F_{f})_{\xi,S,\Phi})_{0}$ for any $\Phi\in\mathcal{S}(X(\mathbb{A}))$ with $\Phi_{\mathbb{f}}={\rm char}_{V^{\prime}(\widehat{\mathbb{Z}})}$. By the above computation we have
@@
Keep the notation being as above. The formal series $F_{f}$ is automorphic cuspidal form on $G(\mathbb{A})$ of weight $l+1$ with respect to G(^)G($\widehat{\mathbb{Z}}$) in the sense of Section 3.
@@
Since $G(\mathbb{Q})=H^{{\rm ss}}(\mathbb{Q})P(\mathbb{Q})$, the claim follows from Lemma 8.2 and Lemma 8.3, and Proposition 8.4. ∎

### v3

@@ 4. Eisenstein series
In this section, we follow [29, Section 4]. For any $l>n+1$ and a Schwartz-function $\Phi_{\mathbb{f}}$ on $V(\mathbb{A}_{\bf f})$, we define a section $f$ of ${\rm Ind}^{G(\mathbb{A})}_{P(\mathbb{A})}\,|\nu|^{s-\frac{n+2}{2}}$ (normalized induction) by
@@
$f_{{\rm fte}}(g_{\mathbb{f}},\Phi_{\mathbb{f}},s)=\int_{{\rm GL}_{1}(\mathbb{A}_{\mathbb{f}})}|t|^{s}\Phi_{\mathbb{f}}(tg_{\mathbb{f}}^{-1}e)dt,\ g_{\mathbb{f}}\in G(\mathbb{A}_{\mathbb{f}})$
and $f_{l,\infty}(g_{\infty},s)$ is the $V_{l}={\rm Sym}^{2l}(\mathbb{C}^{2})$-valued, $K$-equivalent section of ${\rm Ind}^{G(\mathbb{R})}_{P(\mathbb{R})}\,|\nu|^{s-\frac{n+2}{2}}$. See Section 4.2 for a precise definition. Then, we define the associated Eisenstein series as
@@
where $w\in G={\rm SO}(V)$ is defined by $w\cdot e=f$, $w\cdot f=e$, and $w|_{V^{\prime}}={\rm id}_{V^{\prime}}$. Notice that $J(s,\eta,\Phi_{p})$ is also defined for $\eta\in V^{\prime}(\mathbb{Q}_{p})$ with $q(\eta)\neq 0$. Further, by [29, p.633, line -6], $J(s,\eta,\Phi_{p})=\displaystyle\int_{{\rm GL}_{1}(\mathbb{Q}_{p})}\int_{V^{\prime}(\mathbb{Q}_{p})}\psi_{p}((\eta,x))\Phi_{p}(t,tx,-tq^{\prime}(x)))|t|^{s}_{p}\,dxdt$. Since $t$ runs over $\mathbb{Z}^{\times}_{p}$, $\Phi_{p}(t,tx,-tq^{\prime}(x)))$ is preserved under the transformation $x\mapsto x+x_{0}$ for any $x_{0}\in V^{\prime}(\mathbb{Z}_{p})$. By the change of variables, $J(s,\eta,\Phi_{p})=\psi((\eta,x_{0}))J(s,\eta,\Phi_{p})$ for any $x_{0}\in V^{\prime}(\mathbb{Z}_{p})$. Thus, $J(s,\eta,\Phi_{p})=0$ if $\eta\not\in V^{\prime}(\mathbb{Z}_{p})$.
The following lemma is an analogue of [20, p.591, Lemma 3.4] and it is proved by the same argument therein:
@@
Keep the notation as above. For each $g_{p}=m_{p}n_{p}\in P(\mathbb{Q}_{p})=M(\mathbb{Q}_{p})N(\mathbb{Q}_{p})$ with $m_{p}={\rm diag}(t_{p},\ast,t_{p}^{-1})$, it holds
$J(s,\eta,g_{p}\cdot\Phi_{p})=|t_{p}|^{\frac{n+2}{2}}_{p}\psi_{p}(m_{p}n_{p}m^{-1}_{p})J(s,m_{p}^{-1}\eta,\Phi_{p}).$
@@ 4.1. $p$-adic computation: Siegel series
$B_{r,\eta}=\int_{p^{-r}V^{\prime}(\mathbb{Z}_{p})}\psi((\eta,x))\text{char}(p^{r}q(x)\in\mathbb{Z}_{p})\,dx=p^{r(n+2)}\int_{V^{\prime}(\mathbb{Z}_{p})\atop q(x)\equiv 0\,\text{mod $p^{r}$}}\psi\Big(\frac{(\eta,x)}{p^{r}}\Big)\,dx.$
@@ 5. The Weil representation of the Jacobi group $J$
We refer to [13] for the basics of the Weil representation. Let $\psi:\mathbb{Q}\backslash\mathbb{A}\longrightarrow\mathbb{C}^{\times}$ be the standard additive character (cf [19, Section 4.2]). For each $S\in\mathbb{Q}^{\times}$, let $\psi_{S}=\psi(S\ast)=\otimes^{\prime}_{p}\psi_{S,p}$ where we also regard it with an additive character on $Z(\mathbb{A})$ via the identity $Z(\mathbb{A})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{A},\ v(0,0,z)\mapsto z$.
@@ 6.2. Fourier-Jacobi expansions of Eisenstein series
Let $\iota_{{\rm SL}_{2}}=\left(\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right)$ be the Weyl element of ${\rm SL}_{2}$. The following theorem is proved by a standard way (see the proof of [13, p.627, Theorem 3.2]) with the double coset decomposition in Section 6.1. Thus, we omit the proof.
@@ 7.1. Preliminaries at the archimedean place
$\lambda=\sqrt{\beta}\sqrt{2\pi|S|},\ \mu=2\pi\beta\Big(|a|+|S|t^{2}_{n}+|S|(x^{\prime},x^{\prime})\Big).$
@@ 7.2. Computation of Fourier-Jacobi coefficients
Note that (7.16) is a finite sum (see [19, Section 5.2]).
@@
$\phantom{xxxxxxxxxxxx}=n(-\alpha,\alpha x_{\infty},\frac{\alpha}{2}\sigma(x_{\infty},x_{\infty}))v(x_{\infty},0,0){\rm diag}(\sqrt{\beta},\sqrt{\beta}^{-1})\in N(\mathbb{R})M(\mathbb{R})$
@@
$=$ $\varepsilon^{n}_{\infty}\frac{\gamma_{\infty}(1)}{\gamma_{\infty}(S^{n})}\beta^{\frac{n}{4}}\Phi_{\infty}(\sqrt{\beta}x_{\infty})\psi_{\infty}(\tfrac{S\alpha}{2}\sigma(x_{\infty},x_{\infty})).$
@@
(this follows from the argument for [19, Section 5.2] and the proof of [20, Lemma 5.4]).
@@ 8. A uniform proof for the main theorems
In this section, we prove Theorem 1.1 and 1.2. We follow the method in [14], [15], [19] rather than representation theoretic arguments in [18], [20].
@@ 8.1. A formal series
Recall the Laurent polynomial $\widetilde{Q}_{\eta,p}(X_{p})$ depending on $\eta\in V^{\prime}(\mathbb{Q}_{p})$ and $\Phi_{p}={\rm char}_{V^{\prime}(\mathbb{Z}_{p})}$. We write it as $\widetilde{Q}_{\eta,\Phi_{p}}(X_{p}):=\widetilde{Q}_{\eta,p}(X_{p})$ to indicate independence. By Lemma 4.1 and the computation in Section 4, for each $g_{p}=m_{p}n_{p}\in P(\mathbb{Q}_{p})=M(\mathbb{Q}_{p})N(\mathbb{Q}_{p})$ with $m_{p}={\rm diag}(t_{p},\ast,t_{p}^{-1})$, we see that
(8.1) $\widetilde{Q}_{\eta,g^{-1}_{p}\Phi_{p}}(X_{p})=|t_{p}|^{\frac{n+2}{2}}_{p}\psi_{p}(m_{p}n_{p}m^{-1}_{p})\widetilde{Q}_{m_{p}^{-1}\eta,\Phi_{p}}(X_{p})$
@@ 8.2. A compatible family of Eisenstein series
makes up a compatible family of Eisenstein series in the sense of [15, Section 9] for $n$ even and of [14, Section 10] for $n$ odd.
The case when $S>0$ is similarly handled and we also have a compatible family of Eisenstein series for $v=0,-l^{\prime}$. In fact, we have a family of anti-holomorphic Eisenstein series but the above mentioned results also hold in this setting.
@@ 8.3. Automorphy of the formal series
$F_{f}(g):=F(g;\{\alpha_{p}\}_{p})=\sum_{\eta=(a,\xi,S)\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0,\ aS\neq 0}A(\eta;\{\alpha_{p}\}_{p})(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\quad g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A}),$
@@
Now assume $n$ is odd. Let $D$ be a square-free fundamental discriminant. Then for $\eta=(D,D,\overbrace{0,\ldots,0}^{n},\pm 1,\pm 1)$, $q(\eta)=\pm D$, where we choose a sign so that $\pm D>0$. Then by Theorem 4.3, Theorem 4.9, and (8.3), $A(\eta;\{\alpha_{p}\}_{p})=c(|D|)$. By Kohnen-Zagier [23, Theorem 1], $c(|D|)$ is proportional to the central value of $f\otimes\chi_{D}$. The claim follows from [28, p.654]. ∎
@@
Let $\gamma=mn\in P(\mathbb{Q})=M(\mathbb{Q})N(\mathbb{Q})$ with $m={\rm diag}(c,\ast,c^{-1})$. By (8.1) and (3.3), we have
$F_{f}(\gamma g)=\sum_{\eta=(a,\xi,S)\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0,\ aS\neq 0}\psi((\eta,mnm^{-1}))A(\eta;\{\alpha_{p}\}_{p})(mg_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(mg_{\infty})$
$\phantom{xxxxx}=\sum_{\eta=(a,\xi,S)\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0,\ aS\neq 0}|c|^{l-\frac{n}{2}}_{\mathbb{A}_{f}}\cdot|c|^{\frac{n+2}{2}}_{\mathbb{A}_{f}}A(m^{-1}\eta;\{\alpha_{p}\}_{p})(g_{\mathbb{f}})|c|^{(l+1)}_{\infty}\mathcal{W}_{2\pi m^{-1}\eta}(g_{\infty}).$
Clearly, $|c|^{l-\frac{n}{2}}_{\mathbb{A}_{f}}\cdot|c|^{\frac{n+2}{2}}_{\mathbb{A}_{f}}|c|^{(l+1)}_{\infty}=|c|^{l+1}_{\mathbb{A}}=1$ since $c\in\mathbb{Q}^{\times}$. Replacing $m^{-1}\eta$ with $\eta$ does not change the conditions in the summation. Thus, we have the claim. ∎
@@
$(F_{f})_{0}(g)=\sum_{\eta=(a,\xi,\ast)\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0,\ a\ast\neq 0}\Bigg(A(\eta;\{\alpha_{p}\}_{p})(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty})\displaystyle\int_{Z(\mathbb{Q})\backslash Z(\mathbb{A})}\psi(az)dz\Bigg)=0$
[…]
$F_{f}(g)=\displaystyle\sum_{S\in\mathbb{Q}^{\times}}(F_{f})_{S}(g).$
Applying the argument in [13, Proposition 1.3], for each $S\in\mathbb{Q}^{\times}$, we can approximate $(F_{f})_{S}(\widetilde{h}v)$ by $(F_{f})_{S,\Phi}(\widetilde{h}v)\Theta_{\psi_{S}}(\widetilde{h}v)$ where
[…]
and $\Phi$ runs over $\mathcal{S}(X(\mathbb{A}))$. Notice $H^{{\rm ss}}(\mathbb{Q})$ splits under the double covering $\widetilde{H^{{\rm ss}}(\mathbb{A})}\longrightarrow H^{{\rm ss}}(\mathbb{A})$ and $\Theta_{\psi_{S}}(\widetilde{h}v)$ is left $H^{{\rm ss}}(\mathbb{Q})$-invariant. Then, by the argument in [20, Lemma 5.4], for any fixed $v$ ($l\leq v\leq l$), we may check the $H^{{\rm ss}}(\mathbb{Q})$-invariance of the $v$-th component $((F_{f})_{S,\Phi})_{v}$ (defined by (7.19)) of $(F_{f})_{S,\Phi}$ for any $\Phi\in\mathcal{S}(X(\mathbb{A}))$.
@@
Since $F_{f}$ is right G(^)G($\widehat{\mathbb{Z}}$)-invariant by construction and left $P(\mathbb{Q})$-invariant by Lemma 8.3, we may assume $\Phi_{\mathbb{f}}={\rm char}_{V^{\prime}(\widehat{\mathbb{Z}})}$. Assume $S<0$. First, we observe the $l$-th component $((F_{f})_{\xi,S,\Phi})_{l}$ for $\Phi$ whose archimedean component $\Phi_{\infty}$ is Gaussian $\Phi_{S,0}$. Then, applying Section 8.2 and [15, Theorem 10.2] for $n$ even and [14, Theorem 10.5] for $n$ odd, the $l$-th component $((F_{f})_{\xi,S,\Phi})_{l}$ is automorphic, hence left $H^{\rm ss}(\mathbb{Q})$-invariant. It is also corresponding to a holomorphic modular form of weight $l-\frac{n}{2}+1$. Further, $C_{4}(S,\Phi_{\infty})=C_{2}(S,\Phi_{S,0})$ is non-zero by Remark 7.1. Thus, the formal Fourier-Jacobi coefficient
@@
Now we turn to the 0-th component $((F_{f})_{\xi,S,\Phi})_{0}$ for any $\Phi\in\mathcal{S}(X(\mathbb{A}))$ with $\Phi_{\mathbb{f}}={\rm char}_{V^{\prime}(\widehat{\mathbb{Z}})}$. By the above computation we have
@@
Keep the notation being as above. The formal series $F_{f}$ is a cuspidal automorphic form on $G(\mathbb{A})$ of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$) in the sense of Section 3.
@@
Since $G(\mathbb{Q})=H^{{\rm ss}}(\mathbb{Q})P(\mathbb{Q})$, the claim follows from Lemma 8.2 and Lemma 8.3, and Proposition 8.4. The cuspidality follows from Proposition 3.4. ∎